OpenStudy (anonymous):

Determine the slope of the line that contains the points G(8,1) & H(8,-6)

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2}\]

OpenStudy (alexandervonhumboldt2):

substitute points: (1-(-6))/(8-8)=7/0=you canno't divide by 0, thus the line has undefined slope, it is vertical